OpenStudy (anonymous):

2^2x-1=10 i need to solve it in the chunking formula

OpenStudy (earthcitizen):

\[x=2.2\]

OpenStudy (anonymous):

you can simplify this to 4^x = 11. then you can do ln 4^x = ln 11 x ln 4 = ln 11 x = ln 11/ln 4 which gives you 1.729... = x

OpenStudy (anonymous):

where 'ln' is the natural logarithm

OpenStudy (mathmate):

\[2^{2x-1} =10 \ =>\ 2x-1=\frac{log(10)}{log(2)}=> x\approx2.161\] \[2^{2x}-1\ =\ 10 \ => \ 2x=\frac{log(11)}{log(2)} \ =>\ x\ \approx 1.7297\]